### v1

# Harder or Different? Understanding Generalization of Audio Deepfake Detection

Abstract
Recent research has highlighted a key issue in speech deepfake detection: models trained on one set of deepfakes perform poorly on others. The question arises: is this due to the continuously improving quality of Text-to-Speech (TTS) models, i.e., are newer DeepFakes just ‘harder’ to detect? Or, is it because deepfakes generated with one model are fundamentally different to those generated using another model? We answer this question by decomposing the performance gap between in-domain and out-of-domain test data into ‘hardness’ and ‘difference’ components. Experiments performed using ASVspoof databases indicate that the hardness component is practically negligible, with the performance gap being attributed primarily to the difference component. This has direct implications for real-world deepfake detection, highlighting that merely increasing model capacity, the currently-dominant research trend, may not effectively address the generalization challenge.
keywords:
audio deepfake detection, anti-spoofing, generalization
1 Introduction
Recent years have seen tremendous advances in machine learning (ML). One area in which progress has been particularly impressive is that of text-to-speech (TTS) synthesis. It is now possible to generate high-quality, convincing speech signals which mimic closely the voice identity [1, 2, 3]. Numerous online services, e.g. [4, 5, 6], enable this technology to be used by anyone, even those without any technical expertise. While it has plentiful legitimate applications, its accessibility has also given rise to serious threats, e.g. fraud, misinformation, or defamation [7, 8, 9, 10].
A potential solution lies in the ML-driven detection of such deepfakes using, for example, binary classifiers to discriminate between genuine/bonafide and AI-generated speech. The field has witnessed a surge in research, from the development of new detection models [11, 12, 13, 14, 15, 16] to the creation of extensive datasets [17, 18, 19, 20, 21, 22, 23]. Most notably, initiatives such as ASVspoof [24, 25, 26] which were launched to benchmark competing detection solutions, seemingly show impressive progress; lower and lower state-of-the-art error rates are reported on a regular basis [15, 16, 27]. However, reliability in real-world scenarios often remains untested, while there are reports that generalisation to out-of-domain scenarios is wanting. The limited ability to generalize to deepfakes generated using new attack algorithms, or even algorithms that are simply different to those used to create training data, has been and continues to be a source of major concern [17].
Despite generalisation having been a focal point of related research for almost a decade, we remain far from practical deepfake detection solutions which generalise to attacks and acoustic conditions seen in the wild, c.f. Figure 1. We assume that one key to unlocking progress is an understanding of why current detection solutions fail to generalize. The answer can be attributed to two main factors. First, for want of a better term, ‘hardness’: the increasing sophistication of speech deepfakes may make them inherently more challenging to detect. Second, ‘difference’: the characteristics used by a detection model to discriminate between bonafide speech and deepfakes may not generalize across different attack algorithms. This implies that detection failures might not result from a lack of detection model capacity, but from fundamentally different deepfake characteristics and features. Thus, the issue of generalization may be due to either ‘hardness’ or ‘difference’, or a combination thereof.
Contributions. We report a means to decompose the performance discrepancy between training and out-of-domain test data into two components: ‘hardness’ and ‘difference’. We report a study, performed using four different detection models and the ASVspoof 2019 and ASVspoof 2021 datasets, which shows that:
  • When randomly degrading input in the ASVspoof 2019 database, performance is significantly diminished, which we show can be attributed mainly to ‘hardness’.
  • For the ASVspoof 2021 logical access (LA) dataset, which contains variation related to the use of different compression and telephony encoding algorithms, degraded detection performance can be attributed to both ‘hardness’ and ‘difference’.
  • In contrast, for the ASVspoof 2021 deepfake (DF) dataset which contains data collected from other sources, degraded detection performance can be attributed almost exclusively to ‘difference’. This also holds for the In-the-Wild database [17].
These findings suggest that efforts to extend model capacity, while beneficial in the case of in-domain benchmarking, are insufficient and might even be detrimental to the pursuit of generalisable detection solutions. We argue that this calls for the re-focusing of research effort to better understand and address the ‘difference’ between deepfake techniques in the wild.
2 Related Work
A significant research in speech deepfake detection was performed in the scope of ASVspoof challenges [24, 28, 25, 26]. This body of work has established a benchmark that subsequent work has taken as a reference point to assess performance [11, 12, 13, 14, 15, 16]. However, recent observations highlight that promising results obtained using ASVspoof databases do not necessarily translate to reliable detection in real-world scenarios [29, 17]. This finding has prompted the creation of new datasets designed to address the performance gap [23]. Research has also explored the detection of so-called partial-spoofs, where bonafide utterances are segmented and then concatenated with content generated using text-to-speech synthesis [20]. Other directions include the investigation of deepfake detection for singing voices [22]. Despite progress, generalisation [30] issues and robustness [31] remain in focus, new attacks continue to emerge and so the challenge is still as great as ever. Whereas specific studies of generalisation have been reported in other fields, e.g. computer vision and object detection [32, 33], a similar comprehensive analysis of the underlying challenges in speech deepfake detection is lacking. The related work in computer vision serves as inspiration and as a methodological foundation for the study reported in this paper.
3 Methodology
Let $A$ represent a learning rule, encompassing model architecture, hyperparameters, and training configuration. After training on the training partition of database $D$, we evaluate the performance $p$ on the test partition of database $D^{\prime}$. Training and test partitions are always disjoint, but may or may not be sourced from the same database, representing either an in-domain (i.e., $D\to D)$ or out-of-domain (i.e., $D\to D^{\prime})$ evaluation scenario, respectively. Performance is assessed using the Equal Error Rate (EER) which, despite well-known shortcomings, is often used when comparing independent spoofing detection solutions [28, 26].
To gain insights into performance deteriorations stemming from generalisation issues, we compare in-domain performance $p^{A}_{D\rightarrow D}$ to out-of-domain performance $p^{A}_{D\rightarrow D^{\prime}}$ for different detection systems and datasets. To illustrate the idea, we aim to understand why models trained on the data from one database exhibit better detection performance when tested on data from the same database (in-domain) compared to data from a different database (out-of-domain). This performance is expressed in terms of the performance gap
$p^{A}_{D\rightarrow D}-p^{A}_{D\rightarrow D^{\prime}}.$ (1)
which is the difference in detection performance for in-domain and out-of-domain scenarios. It can be decomposed as follows.
$\underbrace{p^{A}_{D\rightarrow D}-p^{A}_{D\rightarrow D^{\prime}}}_{\text{Performance gap}}$ (2)
$=$ $\underbrace{p^{A}_{D\rightarrow D}-p^{A}_{D^{\prime}\rightarrow D^{\prime}}}_{\text{Hardness gap}}+\underbrace{p^{A}_{D^{\prime}\rightarrow D^{\prime}}-p^{A}_{D\rightarrow D^{\prime}}}_{\text{Difference gap}}$
Thus, we break down the performance gap into two components: hardness gap and difference gap. If the performance gap stems primarily from inherently more challenging test data, such as more sophisticated deepfakes, i.e. attacks that produce high-quality speech signals with minimal artefacts, then we anticipate a more significant ‘hardness gap’. On the other hand, if deepfakes contained in the test data exhibit a similar level of artefacts but differ in acoustic conditions compared to the training data, then the performance gap should be attributed to a greater ‘difference’ gap. We expect the same in the case that deepfakes contained in the test data are no more ‘hard’ to detect, if only learning was performed using similar deepfakes among the training data, hence the distinction between what is ‘hard’ and what is ‘different’. If the ”performance gap” is primarily due to ”difference”, then enhancing model capabilities may not address the underlying issue.
Figure 1: Comparison of EER between ASVspoof 2019 (x-Axis) and ASVspoof 2021 (y-axis) for models trained on ASVspoof 2019. None of the model achieves ideal generalization, as denoted by the dashed line.
4 Evaluation
We use four learning models from related work in our experiments: The LCNN model [11], the RawNet2 [34], Whisper-DF [16] and SSL-W2V2 [35]. These models represent the most common approaches to speech deepfake detection, and were state-of-the-art at the time of their publication. When training these models, we use a batch sizes of 16, the Adam optimizer with a learning rate of $10^{-3}$, and allow up to $500$ epochs. To prevent overfitting, we use aggressive early stopping with a delta of $0.5\%$ train EER and a patience of $1$ epoch. Additionally, we remove leading and trailing silence from all data, as models tend to overly rely on non-informative features such as silence [29].
4.1 Databases
We utilize the following databases:
ASVspoof 2019 LA [28]: We use ASVspoof 2019 LA for both training and in-domain testing. It has three disjoint partitions: training, development and evaluation. Importantly, the evaluation partition includes out-of-domain data, namely, attacks not seen during the training and development phases. Since our focus is on using ASVspoof 2019 LA for training and in-domain testing, we disregard these partitions and reorganize the entire dataset into an $80\%$ training and $20\%$ in-domain testing split.
ASVspoof 2021 [36]: To evaluate with out-of-domain data, we utilize samples from the ASVspoof 2021 challenge. This includes the ‘Logical Access’ (LA) dataset and the ‘Deepfake’ (DF) dataset. Compared to ASVspoof 2019, these datasets pose additional challenges. The ASVspoof 2021 LA dataset simulates more realistic channel conditions, with all audio files being processed through real telephony systems, including voice-over-internet-protocol (VoIP) systems and public switched telephone networks (PSTN) [26]. Consequently, detection models must handle compression artifacts, packet losses, and other related issues. On the other hand, the ASVspoof 2021 DF dataset introduces new audio data, specifically, previously unseen deepfakes from the Voice Conversion Challenges of 2018 and 2020 [37, 38]. Additionally, lossy codecs are used for media storage to closely mimic realistic scenarios.
In-the-Wild [17]: This dataset includes both synthesized and genuine audio for 58 politicians and public figures, obtained from social networks and video platforms. It encompasses 20.8 hours of authentic and 17.2 hours of fake audio. This dataset poses a significant challenge, and models trained on ASVspoof 2019 have shown inadequate generalization on it.
4.2 Experiments
Our objective is to assess the decomposition of the performance gap into ‘hardness’ and ‘difference’ components across various scenarios.
Audio Degradation. In our initial experiment, we deliberately lower the quality of the test data by aggressively truncating input length. In this scenario, with respect to Equation 2, $D$ denotes the original ASVspoof 2019 dataset with $8$ seconds (s) of input, while $D^{\prime}$ represents the same data where each sample is randomly truncated to $0.25s$. For models that require a minimum audio length (such as WhisperDF and SSL-W2V2), the truncated input is looped. Detailed outcomes are provided in Section 5.1.
Unseen databases. The second, third, and fourth experiments evaluate the performance gap between ASVspoof 2019 LA and new databases: ASVspoof 2021 LA, DF, and In-the-Wild, respectively. These scenarios represent progressively increasing complexity levels. While ASVspoof 2021 LA includes the same attacks [^1] In this context, an ‘attack’ refers to a distinct Text-to-Speech (TTS) synthesis algorithm. Introducing a new and unseen attack, therefore, means presenting a new and previously unencountered category of audio spoof or audio deepfake to the model. as ASVspoof 2019 but incorporates lossy compression and telephone encodings, ASVspoof 2021 DF introduces partially new attacks (sourced from the voice conversion challenge databases [37, 38]). Lastly, the In-the-Wild dataset consists entirely of new, unseen, and potentially high-quality attacks due to its recency compared to ASVspoof 2019. The results of these experiments are presented in Section 5.2 and Section 5.3.
5 Results
5.1 Audio Degradation
Table 1: Decomposition of the performance gap into ‘hardness’ and ‘difference’ for the ASVspoof 2019 dataset, when input length is truncated for test data. A significant portion of the performance gap can be attributed to ‘hardness’ (blue). Surprisingly however, for some models, there also is a rather significant contribution from ‘difference’ (red). Results presented as mean EER ± standard deviation across five individual trials.
Model Input (s) Perform. Hardn. Diff.
Train Test Gap Gap Gap
LCNN 8.0 0.25 $17.7\pm 0$ $13.4\pm 2$ $4.2\pm 2$
RawNet2 8.0 0.25 $33.4\pm 4$ $30.4\pm 3$ $3.0\pm 5$
SSL W2V2 8.0 0.25 $23.1\pm 2$ $11.8\pm 2$ $11.2\pm 3$
WhisperDF 8.0 0.25 $25.1\pm 1$ $8.4\pm 1$ $16.6\pm 1$
Recall that as per Equation 2, the ‘performance gap’ is the sum of the ‘hardness gap’ and ‘difference gap’. Table 1 illustrates this breakdown of performance degradation due to reduced input length. We observe that the main reason for this is the ‘hardness gap’, as marked in blue. As expected, diminishing the amount of information available to the model inherently increases the problem’s ‘hardness’. The RawNet2 model experiences the most pronounced effect: a $33.4\%$ increase in EER, of which $30.4\%$ EER can be attributed to increased ‘hardness’, while only $3.0\%$ EER accord to the ‘difference’ in conditions. Similarly, the LCNN model’s performance deteriorates by $17\%$ EER, of which $13\%$ EER pertain to an increase in ‘hardness’. Larger, pre-trained models like SSL W2V2 and WhisperDF, exhibit a less pronounced sensitivity to the ‘hardness gap’, but are still significantly affected. Interestingly, there also is a significant contribution to ‘difference’. This is discussed in section 6.
5.2 ASVspoof 2021 LA
Table 2 details the results for the ASVspoof 2021 LA dataset, which contains the same audio files as ASVspoof 2019, but which have been altered significantly through the use of encoding and compression techniques. Surprisingly, contrary to expectations these modifications would increase the dataset’s difficulty, this findings reveal a different scenario. Although performance deterioration is less pronounced than in the cases of input data reduction, the primary factor for the performance gap is attributed to ‘difference’. Notably, the LCNN model, the smallest among the four analyzed, shows a blend of both hardness and difference in its performance decline. In contrast, larger models predominantly link their diminished performance to ‘difference’. We will further discuss the implication of this observation on model generalization in Section 6.
Table 2: Decomposition of the performance gap when training on ASVspoof 2019 and evaluating on ASVspoof 2021 LA. For most models, the performance gap can be attributed solely to ‘difference’ (red).
Model Performance Hardness Difference
Gap Gap Gap
LCNN $20.5\pm 2$ $7.2\pm 3$ $13.4\pm 3$
RawNet2 $15.3\pm 7$ $-0.6\pm 7$ $15.9\pm 1$
SSL W2V2 $11.8\pm 2$ $0.6\pm 0$ $11.2\pm 2$
WhisperDF $10.7\pm 2$ $0.8\pm 0$ $9.9\pm 2$
5.3 ASVspoof 2021 DF and In-The-Wild
Table 3: Decomposition of performance gap between ASVspoof 2019 and ASVspoof 2021 DF. The performance gap can be attributed solely to ‘difference’ (red).
Model Performance Hardness Difference
Gap Gap Gap
LCNN $26.6\pm 3$ $1.7\pm 2$ $25.0\pm 3$
RawNet2 $18.1\pm 7$ $-2.2\pm 7$ $20.3\pm 2$
SSL W2V2 $14.2\pm 3$ $0.1\pm 0$ $14.0\pm 3$
WhisperDF $15.2\pm 2$ $0.1\pm 0$ $15.1\pm 2$
Table 4: Decomposition of performance gap between ASVspoof 2019 and In-the-Wild. The performance gap can be attributed solely to ‘difference’ (red).
Model Performance Hardness Difference
Gap Gap Gap
LCNN $78.2\pm 15$ $-1.5\pm 2$ $79.7\pm 15$
RawNet2 $40.6\pm 7$ $-6.4\pm 7$ $47.0\pm 3$
SSL W2V2 $30.3\pm 5$ $0.1\pm 0$ $30.2\pm 5$
WhiserDF $31.4\pm 5$ $-0.5\pm 0$ $31.9\pm 5$
This final experiment assesses the trade-off between hardness and difference new, unseen attacks. For this purpose, ASVspoof 2021 DF, containing partially unseen attacks, and In-the-Wild, encompassing completely unseen and more recent attacks, are utilized. Table 3 displays the results for ASVspoof 2021 DF, where an average performance gap of $14\%$ to $26\%$ in EER is observed, predominantly attributed to the ‘difference gap’. The evaluation on In-the-Wild, c.f. Table 4, reveals an even more substantial performance gap ($31\%$ to $78\%$), which is entirely due to the ‘difference gap’. Consequently, all models perceive the new attacks as a domain shift, characterized not by an escalation in difficulty or complexity but rather as encountering new, unseen, and distinct data and/or features.
6 Discussion and Implications
These results are astounding. First, we had expected data augmentation to increase ‘hardness’. However, for input length truncation and especially for compression and telephone encodings in ASVspoof 2021 LA, we observe a marked increase in ‘difference’, c.f. Section 5.2. To us, this is very surprising, since introducing more noise to the test data could be considered the archetype of increasing data difficulty. Second, whenever models are presented with new data, i.e. unseen attacks, the lack of generalization performance is attributed near-exclusively to ‘difference’. This is in stark contrast to similar experiments in the vision domain, where Lu et Al. evaluate the generalization of models trained on CIFAR-10 to the out-of-domain database CIFAR-10.2 [33]. Here, a ‘performance gap’ of about $15\%$ accuracy is, on average, decomposed into a ‘difference gap’ of $5\%$ and a ‘hardness gap’ of $10\%$. This more balanced decomposition is in stark contrast to our experiments on ASVspoof 2021 DF and In-the-Wild, where the ‘difference gap’ is the sole contributor to performance degradation.
Addressing the challenge of ‘harder’, data could potentially be within reach by developing more sophisticated models [33]. However, ‘difference’ poses a more elusive challenge and may not be effectively addressed by merely increasing model complexity. Recall that machine learning fundamentally is pattern recognition. Patterns that fall outside the training data are, by definition, difficult for a model to comprehend, given that it derives all its knowledge from the statistical aggregation of input-output relationships present in the training data. Thus, the consistent prevalence of the ‘difference’ gap across all experiments, especially as the dominant factor in ASVspoof 2021 and In-the-Wild, might explain why achieving true out-of-domain generalization has been so problematic in previous research. It appears that deepfake-detection and anti-spoofing models are overly specific to their training data, with even minor deviations treated as outside the training data distribution (i.e. ‘different’).
We hypothesize that models learn specific features that perform well for the current database, but which are so specific that they may not necessarily transfer well to other out-domain databases. An example from related work is the ‘silence-shortcut’ in the ASVspoof 2019 dataset [29]: Here, the models can use the length of leading silence to achieve near-perfect performance. The length of the silence correlates with the target variable in ASVspoof 2019 train; and even generalizes to the ASVspoof 2019 test data, but is obviously not a semantically meaningful feature that can be expected to work in the real world. It may be likely that there are other such ‘shortcuts’ [39] in voice anti-spoofing datasets, which may be harder for humans to identify, but which nevertheless are very predictive clues that are exploited by anti-spoofing models. The existence of such ‘shortcuts’ would align with our observation of the dominance of the ‘difference gap’, which is not nearly as pronounced in, for example, the vision domain [33].
Therefore, while increasing model size by adding more parameters may prove beneficial in other domains, such as generating text with large language models, it might not be as effective in the realm of supervised classification for audio anti-spoofing and deepfake detection. At the same time, the use of self-supervised learning (SSL), exemplified by SSL-W2V2 and WhisperDF, appears to be a promising approach, as indicated by the overall smaller ‘performance’ gaps. This might be because such pre-trained models, when employed for voice anti-spoofing, usually are not fine-tuned. This could allow them to extract less specific, and thereby better generalizing features.
7 Future Work and Conclusion
This paper presents a study demonstrating that the performance gap in anti-spoofing generalization can be dissected into a ‘hardness’ and a ‘difference’ component. Our hypothesis is that current failures in generalization are predominantly due to the significant yet inadequately addressed influence of the ‘difference’ gap, suggesting that detection models overfit to the training distribution.
Future research should aim to deepen the understanding of this issue by systematically analyzing and breaking down errors across a wide range of audio augmentations (such as, potentially, augmentation by band-pass filters, compression, room impulse responses, Gaussian noise, time- and pitch-shifting, etc.), as well as between different attacks, e.g. between autoregressive and transformer-based TTS models. Such investigations could provide a more comprehensive view of the conditions under which models achieve generalization versus those in which they latch onto features highly specific to the training dataset. Such research may be more beneficial than current approaches, which simply increase model size, a strategy which may address only the ‘hardness’, but ignore the key ‘difference’ problem of voice authenticity classification.